Calculate for each circuit
a) Ct
b)Qt
c) Energy in the circuit
CALCULAR DE CADA CIRCUITO a. CT c. ENERGIA EN EL CIRCUITO เเเ.
Answer #1

a) The total capacitance is

C_T=(1/6+1/(4+2))^{-1}=(2/6)^{-1}=3,mu F

b) The total charge in the circuit is

Q_T=C_TV=3 imes 12=36,mu C

c) The total energy is

- 2 ET-CA/2 = 3 x 122/2 = 216p5 = 2.16 x 10-4J